Circular No. 6770

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1997 V1 (LARSEN) Following are preliminary parabolic orbital elements from 35 observa-

tions, Nov. 3–5; it is possible that the comet is of short period.

T = 1997 Feb. 1.11 TT  ω = 60.^{o}16
Ω = 227.80 }2000.0
q = 1.9165 AU  i = 14.60
1997TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Oct. 29 0^{h}15.^{m}49 +13^{o}03.^{′}3 2.746 3.649 151.^{o}3 7.^{o}5 15.8
Nov. 3 0 14.25 +12 30.5 2.826 3.693 146.3  8.6 15.9
8 0 13.40 +12 00.2 2.913 3.737 141.2  9.6 16.0
13 0 12.97 +11 32.7 3.005 3.781 136.1 10.5 16.2
18 0 12.94 +11 08.4 3.102 3.825 131.1 11.2 16.3
23 0 13.30 +10 47.1 3.203 3.869 126.1 11.9 16.4
28 0 14.04 +10 29.1 3.309 3.912 121.2 12.5 16.5

SUPERNOVA 1997dq IN NGC 3810 S. Nakano, Sumoto, Japan, reports the discovery by Masakatsu Aoki,

Tsukioka-cho, Toyama, of an apparent supernova (mag 15.0) on seven unfiltered CCD frames taken on Nov. 2 (limiting mag ∼ 18.5) and confirmed on three frames taken on Nov. 3, using a 0.43-m f/6 reflector. The supernova is located at α = 11^{h}40^{m}55^{s}.90, δ = +11^{o}28′45′′.7 (equinox 2000.0), which is 43′′ west and 29′′ north of the center of NGC 3810. No star is present at this location on his numerous previous frames, taken between 1996 Dec. 21 and 1997 May 17.

S. Jha, P. Challis, P. Garnavich, and R. Kirshner, Harvard-Smithson-ian Center for Astrophysics, report that a spectrum of SN 1997dq, taken by Challis and Jha at the Multiple Mirror Telescope on Nov. 5.5 UT, indicates that it is a type-Ib supernova near maximum light. The spectrum exhibits features of Fe II, He I, C II, O I, and Ca II. The recession velocity of the host galaxy, measured from narrow Hα emission, is 900 km/s.

COMET 22P/KOPFF S. Nakano, Sumoto, Japan, reports that CCD observations by T. Oribe

(Saji Observatory) and A. Sugie (Dynic Observatory) on Nov. 2.72 and 3.72 UT show that this comet is brighter than expected, at m_{1} = 16.9–17.0.

1997 November 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green